OpenStudy (anonymous):

x^3+4x^2+8x+32 factor four term polynomial by grouping

OpenStudy (noelgreco):

Take the first two term, and factor out the GCF. Do the same with the third and fourth terms. You should see the answer in the result.

OpenStudy (anonymous):

= x^2(x+4)-8(x+4) =(x^2-8)(x+4)

OpenStudy (anonymous):

..?